Changes between Version 313 and Version 314 of docs/CommissioningPlan


Ignore:
Timestamp:
01/12/23 18:37:29 (2 years ago)
Author:
kradhakrishnan
Comment:

--

Legend:

Unmodified
Added
Removed
Modified
  • docs/CommissioningPlan

    v313 v314  
    180180||P2.5||[wiki:/docs/CommissioningPlan/P2.5 TCS communication test]||1||C3||[[span(style=color: green, complete )]]||
    181181||P2.8||[wiki:/docs/CommissioningPlan/P2.8 Verification of the mask centering algorithm and final contrast with SOUL-AO and turbulence at ASM]||3||C3||[[span(style=color: green, complete (analysis pending))]]||
    182 ||P2.9||[wiki:/docs/CommissioningPlan/P2.9 Star center Fourier mode images]||1||C3||[[span(style=color: blue, Scheduled (with remote support)  )]]||
     182||P2.9||[wiki:/docs/CommissioningPlan/P2.9 Star center Fourier mode images]||1||C3||[[span(style=color: green, complete (remotely)  )]]||
    183183||P2.10||[wiki:/docs/CommissioningPlan/P2.10 Test of all the remaining templates for sky observations]||2||C3||[[span(style=color: green, complete (rest of the testing on-sky) )]]||
    184184|| || Total without contingency||12 days  ||
     
    203203**__Com-Run-1__**
    204204||'''#'''||'''Activity Name / Observing Case'''||'''Duration (hours)'''||'''Seeing Conditions'''||'''Category'''||'''Status'''||
    205 || P3.1 ||[wiki:/docs/CommissioningPlan/P3.1 Telescope control – sending preset, telescope mode, offset, and focal plane geometry estimation]||3||medium / bad||C5||[[span(style=color: blue, ongoing  )]]||
     205|| P3.1 ||[wiki:/docs/CommissioningPlan/P3.1 Telescope control – sending preset, telescope mode, offset, and focal plane geometry estimation]||3||medium / bad||C5||[[span(style=color: green, complete  )]]||
    206206|| P3.2 ||[wiki:/docs/CommissioningPlan/P3.2 SOUL-AO verification]||6||good / medium||C5 + C7||[[span(style=color: blue, yet to start  )]]||
    207207|| P3.3 ||[wiki:/docs/CommissioningPlan/P3.3 Pupil alignment] (for checking the pupil position and working of the script)||2||bad||C5 + C6||[[span(style=color: green, complete  )]]||
    208208|| P3.4 ||[wiki:/docs/CommissioningPlan/P3.4 PSF Alignment]||3||good / medium||C5 + C6||[[span(style=color: green, complete  )]]||
    209 || P3.5 ||[wiki:/docs/CommissioningPlan/P3.5 Verify the ADC rotation on sky]||4||medium / bad||C5||[[span(style=color: blue, ongoing  )]]||
     209|| P3.5 ||[wiki:/docs/CommissioningPlan/P3.5 Verify the ADC rotation on sky]||4||medium / bad||C5||[[span(style=color: green, complete  )]]||
    210210|| P3.6 ||[wiki:/docs/CommissioningPlan/P3.6 Checking field-stabilized mode/bearing rotation service]||2||bad||C5||[[span(style=color: blue, ongoing  )]]||
    211211|| P3.7 ||[wiki:/docs/CommissioningPlan/P3.7 Internal TT loop working and performance] (in different magnitude)||3||good / medium||C5 + C7||[[span(style=color: blue, ongoing )]]||
     
    222222|| || **Total duration with 15% contingency (rounded up)**||**7** nights||
    223223
    224 [[span(style=color: blue, **Com-Run-1 (6-11 January 2023) is ongoing**)]]. The Com-Run-1 activity log in this [https://docs.google.com/document/d/1Rj-zw9aI5EFbsafOKYUUP225DHgLRJv6JivtNxakLho/edit# link].
     224[[span(style=color: green, **Com-Run-1 (6-11 January 2023) is complete**)]]. The Com-Run-1 activity log in this [https://docs.google.com/document/d/1Rj-zw9aI5EFbsafOKYUUP225DHgLRJv6JivtNxakLho/edit# link].
    225225
    226226'' Phase 4 science verification of the observing cases (1-4) may start after Com-Run-1. ''